# Temporal reasoning for timeline summarisation in social media

### v2

@@ A.2 Experiment on teacher model
After fine-tuning the teacher model, we used the TEMPREASON dataset tan2023towards to evaluate its performance pre- and post-fine-tuning. Specifically, we focused on the L3 part of the dataset, which deals with event-event relations, to determine whether the model could accurately infer the temporal sequence between two events. We use F1 as the evaluation metrics, the experimental results show that the fine-tuned model achieved a 0.07 improvement in this metric compared to the pre-fine-tuning model, highlighting its enhanced ability to infer event-event temporal relations.
@@ A.3 Joint learning
When fine-tuning Phi using joint learning, we mix the timeline summarization and temporal reasoning datasets and shuffle their order. We then fine-tune Phi using this mixed dataset. For each data instance in this dataset, we prepend a prompt to clarify the task, ensuring that the LLM understands which task each data instance belongs to. For example, for timeline summarization data, we add the prompt "This is a timeline summarization task, your task is to summarize the provided timeline + timeline summarisation text". Similarly, for temporal reasoning tasks, we would prepend a different prompt to clarify the task at hand, ensuring that Phi can learn to apply its knowledge accordingly to each type of task.

### v3

@@ 6 Conclusions
Acknowledgments
This work was supported by a UKRI/EPSRC Turing AI Fellowship to Maria Liakata (grant ref EP/V030302/1) and the Alan Turing Institute (grant ref EP/N510129/1). This work was also supported by the Engineering and Physical Sciences Research Council [grant number EP/Y009800/1], through funding from Responsible Ai UK (KP0016) as a Keystone project lead by Maria Liakata.
@@ A.2 Experiment on teacher model
After fine-tuning the teacher model, we used the TEMPREASON dataset Tan et al. (2023) to evaluate its performance pre- and post-fine-tuning. Specifically, we focused on the L3 part of the dataset, which deals with event-event relations, to determine whether the model could accurately infer the temporal sequence between two events. We use F1 as the evaluation metrics, the experimental results show that the fine-tuned model achieved a 0.07 improvement in this metric compared to the pre-fine-tuning model, highlighting its enhanced ability to infer event-event temporal relations.
@@ A.3 Joint learning
[…]